# Phase-shifted remote photoplethysmography for estimating heart rate and blood pressure from facial video

### v3

@@ 4.3 Experimental results on V4V database
Table 4 presents the results of estimating blood pressure on the V4V database. Our proposed algorithm achieved MAE and RMSE of 13.64 and 16.78 for estimating SBP and of 9.40 and 11.90 for estimating DBP, respectively. There are very few literatures which reported the performance for estimating blood pressure on the V4V database. We compared the performance of the proposed method to previous algorithms proposed by Schrumpf et al. [2021] and Hamoud et al. [2023]. As shown in Table 4, our BBP-Net achieved lower MAE and RMSE in estimating both SBP and DBP, with significant margins.
4.4 Ablation study
Table 5: Comparative study using various window lengths for estimating heart rate on the MMSE-HR database. The error rate is measured based on MAE.
Window (sec)
@@ 4.4 Ablation study
Ablation study was conducted to analyze the effects heuristic parameters and components of proposed method. Table 5 present the results of estimating heart rate on the MMSE-HR database using different window lengths. The window length of image sequences is an important heuristic parameter, which directly related with computational power and latency of deep learning models. While small window length is advantageous for reducing latency, it causes loss of temporal information, which can lead to decreased accuracy in estimating heart rate. As shown in Table 5, MAE for estimating heart rate is improved as the window length increases in both cases of using facial and acral rPPG signals. When the window length was set to 2 seconds, the MAE was significantly increased because the input sequence length is smaller than the receptive field of the DRP-Net. In experiments, we selected the window length of 6 seconds to achieve lower error rates with reduced latency compared with previous methods.
Table 6: The effectiveness of data augmentation and $L_{pv}$.
@@
Table 6 presents ablation study on the MMSE-HR database for demonstrating the effectiveness of data augmentation and our time domain loss $L_{pv}$. Applying data augmentation reduced 10.05% of MAE for estimating heart rate from facial rPPG signals from 1.99 BPM to 1.79 BPM, simultaneously reducing 15.98% of RMSE. This result demonstrates that augmentation of bradycardia and tachycardia data is beneficial for reducing outlier predictions of low and high heart rates. Additionally, when applying $L_{pv}$, MAE of the DRP-Net for estimating heart rate from facial rPPG signals was reduced by 0.56% from 1.79 BPM to 1.78 BPM, and RMSE was reduced by 5.53% from 4.52 to 4.27. Experimental results in Table 6 demonstrate that both data augmentation and $L_{pv}$ are advantageous for improving the accuracy of estimated heart rate.
Fig. 7 presents predicted heart rate computed from facial rPPG signals and their reference heart rate. Fig. 7(a) to Fig. 7(e) present the results on the five folds of the MMSE-HR database, and Fig. 7(f) presents the result on the V4V database. The predicted and reference heart rates are almost linearly correlated on the MMSE-HR database, and Pearson correlation r is larger than 0.90 on the entire folds. The linear correlation between the predicted and reference heart rates shows that our proposed method is accurate over the entire range of heart rate. While Pearson correlation is 0.75 on the V4V database, our proposed method shows lower MAE and RMSE compared to previous methods as presented in Table 3.
Figure 7: Predicted heart rate computed from facial rPPG signals and their reference heart rate. Pearson correlation between the predicted and reference heart rate is denoted as r.
Table 7: The effectiveness of phase-shifted rPPG signals and scaled sigmoid function for estimating blood pressure.
Facial rPPG Acral rPPG Scaled sigmoid SBP DBP
function MAE (mmHg) RMSE MAE (mmHg) RMSE
@@
To analyze the performance of BBP-Net for estimating blood pressure, ablation study was conducted to demonstrate the effectiveness of utilizing dual rPPG signals and scaled sigmoid function. Table 7 presents MAE and RMSE for estimating SBP and DBP from different types of rPPG signals. While the MAEs for estimating SBP from facial and acral rPPG signals were 12.51 mmHg and 13.18 mmHg, it was reduced to 10.19 mmHg when utilizing both rPPG signals. This result indicates that the temporal discrepancy in pulse waves at different physiological sites contributes to reducing the error for estimating blood pressure. Table 7 shows the performance of blood pressure estimation with and without the scaled sigmoid function in BBP-Net. Without the scaled sigmoid function, the MAEs for SBP and DBP were increased by 9.38 mmHg and 3.06 mmHg, respectively. These experimental results demonstrate that constraining predicted blood pressure into a bounded range is effective to reduce the error by eliminating outlier estimates.
4.5 International standard
@@
We analyze the performance of the proposed method for estimating blood pressure based on the international standard of the British Hypertension Society (BHS) [O’Brien et al., 1990]. The BHS standard evaluates the percentages of estimates which satisfy the MAEs lower than 5 mmHg, 10 mmHg, and 15 mmHg. For example, to obtain Grade A of the BHS standard, more than 60%, 85%, and 95% of estimates should satisfy the MAEs lower than 5 mmHg, 10 mmHg, and 15 mmHg, respectively. Table 8 presents the percentages within three thresholds for MAE for estimating SBP and DBP on the MMSE-HR and V4V databases. Our proposed method achieved Grade B on the MMSE-HR database and Grade C on the V4V database for estimating DBP.
@@
Table 8: The criteria of the BHS standard and the accuracy of estimating blood pressure based on three thresholds for MAE.
@@
5 Conclusion
This paper proposes a 2-stage deep learning framework consisting of DRP-Net and BBP-Net for estimating heart rate and blood pressure from facial videos. In this paper, we introduce the concept of phase-shifted rPPG signals for analyzing phase discrepancy of pulse waves at facial and acral regions. The DRP-Net extracts facial and acral rPPG signals based on 3D convolution and Siamese-structured heads, and a time domain loss is proposed to supervise the scale of the estimated rPPG signals. The BBP-Net estimates SBP and DBP values within a bounded range from the phase-shifted rPPG signals. To reduce the error, we employed a data augmentation method based on a frame interpolation model. Experiments were conducted on the MMSE-HR and V4V databases, and our proposed method achieved superior performance with a significant margin compared to previous methods. Ablation study was thoroughly performed to demonstrate the effectiveness of the phase-shifted rPPG signals for estimating heart rate and blood pressure.
Acknowledgements
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation(IITP)-Innovative Human Resource Development for Local Intellectualization program grant funded by the Korea government(MSIT)(IITP-2024-RS-2024-00439292)

### v4

@@ IV-C Experimental results on V4V database
IV-D Ablation study
TABLE VI: Comparative study using various window lengths for estimating heart rate on the MMSE-HR database. The error rate is measured based on MAE.
Window (s)
@@ IV-D Ablation study
Ablation study was conducted to analyze the effects of the heuristic parameters and components of the proposed method. Table VI presents the results of heart rate estimation on the MMSE-HR database using different window lengths, numerically verifying the trade-off between efficiency and informativeness. The window length of image sequences is an important heuristic parameter, which directly related with computational power and latency of deep learning models. A longer window length increases the time required to collect input image sequences, preprocessing duration, and the model’s computational cost. While small window length is advantageous for reducing latency, it causes loss of temporal information, which can lead to decreased accuracy in estimating heart rate. As shown in Table VI, MAE for estimating heart rate is improved as the window length increases in both cases of using facial and acral rPPG signals. However, in the case of facial rPPG, the reduction in heart rate estimation error is more significant. These results suggest that with longer temporal information, focusing on learning the periodic patterns using $L_{freq}$ is beneficial for heart rate estimation. When the window length was set to 2 seconds, the MAE was significantly increased because the input sequence length is smaller than the receptive field of the DRP-Net. In experiments, we selected the window length of 6 seconds to achieve lower error rates with reduced latency compared with previous methods.
Table VII presents the ablation study on the MMSE-HR database to demonstrate the effectiveness of proposed training methods, such as data augmentation and time-domain loss $L_{pv}$. Applying data augmentation reduced 10.05% of MAE for estimating heart rate from facial rPPG signals from 1.99 BPM to 1.79 BPM, simultaneously reducing 15.98% of RMSE. This result demonstrates that augmentation of bradycardia and tachycardia data is beneficial for reducing outlier predictions of low and high heart rates. Additionally, when applying $L_{pv}$, MAE of the DRP-Net for estimating heart rate from facial rPPG signals was reduced by 0.56% from 1.79 BPM to 1.78 BPM, and RMSE was reduced by 5.53% from 4.52 to 4.27. Experimental results in Table VII demonstrate that both data augmentation and $L_{pv}$ are advantageous for improving the accuracy of estimated heart rate.
TABLE VII: The effectiveness of data augmentation and $L_{pv}$.
@@
Fig. 7 presents Pearson correlation and the Bland-Altman plot of the predicted and reference heart rate. Fig. 7(a) to Fig. 7(e) and Fig. 7(g) to Fig. 7(k) presents the results from the five folds of the MMSE-HR database. Fig. 7(f) and Fig. 7(l) presents the result from the V4V database. The predicted and reference heart rates exhibit an almost linear correlation on the MMSE-HR database, with the Pearson correlation coefficient (r) exceeding 0.90 across all folds. The linear correlation between the predicted and reference heart rates shows that our proposed method is accurate over the entire range of heart rate. The proposed data augmentation method improved the estimation of both tachycardia and bradycardia data. However, bradycardia in fold 4 shows higher estimation errors due to the significant gap between the distributions of the original training and test datasets. While Pearson correlation is 0.75 on the V4V database, our proposed method shows lower MAE and RMSE compared to previous methods as presented in Table IV. The Bland-Altman plots illustrate a uniform distribution of positive and negative errors. The unbiased trend suggests strong generalizability of our proposed model.
Figure 7: Predicted heart rate computed from facial rPPG signals and their reference heart rate. In subfigures (a)–(f), Pearson correlation between the predicted and reference heart rate is denoted as r. Subfigures (g)–(l) show the Bland-Altman plot of predicted and reference heart rate. The vertical and horizontal axes represent the signed error between predicted and reference heart rate and the reference HR values, respectively.
TABLE VIII: The effectiveness of phase-shifted rPPG signals and scaled sigmoid function for estimating blood pressure.
Facial rPPG Acral rPPG Scaled sigmoid function SBP DBP
MAE (mmHg) RMSE MAE (mmHg) RMSE
@@
To analyze the performance of BBP-Net for estimating blood pressure, ablation study was conducted to demonstrate the effectiveness of utilizing phase-shifted rPPG signals and scaled sigmoid function. Table VIII presents MAE and RMSE for estimating SBP and DBP from different types of rPPG signals. While the MAEs for estimating SBP from facial and acral rPPG signals were 12.51 mmHg and 13.18 mmHg, it was reduced to 10.19 mmHg when utilizing both rPPG signals. This result indicates that the temporal discrepancy in pulse waves at different physiological sites contributes to reducing the error for estimating blood pressure. Table VIII shows the performance of blood pressure estimation with and without the scaled sigmoid function in BBP-Net. Without the scaled sigmoid function, the MAEs for SBP and DBP were increased by 9.38 mmHg and 3.06 mmHg, respectively. These experimental results demonstrate that constraining predicted blood pressure into a bounded range is effective to reduce the error by eliminating outlier estimates.
IV-E Cross skin tone testing
Figure 8: Cross skin tone testing. The horizontal axis represents the Fitzpatrick skin type, while the blue, green, red, and yellow bars represent the MAEs of HR from facial rPPG, HR from acral rPPG, SBP, and DBP, respectively.
Cross skin tone testing was conducted to evaluate the robustness of the proposed DRP-Net and BBP-Net across various skin tones. Skin tones were divided into four folds according to the Fitzpatrick skin type [56], which were type I-II, type III, type IV, and type V-VI. Fig. 8 presents the cross-validation results by skin tone in the form of bar graphs, where the left and right axes represent heart rate and blood pressure errors in MAE, respectively. The acral rPPG-based heart rate estimation error was observed to be 0.77 BPM for lighter tones and 6.03 BPM for darker tones, indicating that it is more challenging to detect periodic signals as skin tone becomes darker. Similarly, the average SBP and DBP estimation errors were highest for skin types V-VI. These findings suggest that the robustness of the proposed pipeline appears to decrease for darker skin tones, indicating a limitation that requires further investigation in future research.
IV-F Computational complexity
TABLE IX: Comparison of computational cost.
Methods Parameters MACs
PhysNet [28] 0.73 M 65.19 G
TS-CAN [57] 3.91 M 61.96 G
AutoHR [41] 0.99 M 189.22 G
EfficientPhys-C [48] 3.84 M 31.32 G
PhysFormer [12] 7.03 M 47.01 G
PhysFormer++ [29] 9.79 M 49.85 G
Ours 0.74 M 38.11 G
Table IX presents the number of parameters and the multiply-accumulates (MACs) for both the previous and proposed heart rate estimation models. The proposed DRP-Net achieves the lowest complexity and heart rate estimation error compared to state-of-the-art methods. Notably, transformer-based models [14, 31] have more than 10 times the number of parameters compared to the proposed model. While PhysNet [30] has a similar number of parameters, it requires approximately 1.71 times more MACs. Additionally, the complexity of a blood pressure estimation model, which shows 0.85 M parameters and 15.6 M MACs. Moreover, the parameters and MACs of the blood pressure estimation model are 0.85 M and 15.6 M, respectively. The efficient 2-stage deep learning framework allows for the acquisition of multiple physiological information with minimal computational cost.
IV-G International standard
@@
We analyze the performance of the proposed method for estimating blood pressure based on the international standard of the British Hypertension Society (BHS) [58]. The BHS standard evaluates the percentages of estimates which satisfy the MAEs lower than 5 mmHg, 10 mmHg, and 15 mmHg. For example, to obtain Grade A of the BHS standard, more than 60%, 85%, and 95% of estimates should satisfy the MAEs lower than 5 mmHg, 10 mmHg, and 15 mmHg, respectively. Table X presents the percentages within three thresholds for MAE for estimating SBP and DBP on the MMSE-HR and V4V databases. Our proposed method achieved Grade B on the MMSE-HR database and Grade C on the V4V database for estimating DBP.
@@
TABLE X: The criteria of the BHS standard and the accuracy of estimating blood pressure based on three thresholds for MAE.
@@
V Discussion
In this paper, we aim to discover direct clues for blood pressure estimation by extracting phase-shifted rPPG signals from facial videos. We demonstrate the superiority of the proposed deep learning framework and detailed training techniques through comparative experiments with previous methods and ablation studies. However, validating the phases of facial rPPG signals is challenging due to the absence of ground truth PPG signals measured from the face. The temporal discrepancy between facial rPPG and acral rPPG is a crucial factor in improving blood pressure estimation performance, and insufficient validation could be considered a limitation of this study.
To address this limitation, we visualized the spatial and temporal attention of DRP-Net to identify the spatiotemporal locations where the model assigned higher weights in the frames. In Fig. 9, the temporal attention exhibited a periodic pattern similar to the physiological signals, particularly with a phase closely aligned with facial rPPG. The spatial attention was designed with a size of $4\times 4$, as the spatial dimension of the feature map is reduced. Fig. 10 illustrates the visualization of spatial attention, overlaid on the clip-averaged images. The visualization results show that higher scores were assigned to facial regions across various skin tones. Notably, for the motion data in the second row, the score is also higher in regions where skin is mainly present.
These attention visualization results suggest that the model has learned the periodicity of emphasized subtle skin tone changes in each facial frame. However, to reliably validate facial rPPG signals, it is necessary to measure facial PPG signals utilizing sensors attached to the face. Future work will focus on constructing a real dataset that includes PPG signals from various body sites, with two main objectives. First, verifying rPPG signals extracted from captured body sites using deep learning models. Second, developing a model for blood pressure estimation based on the temporal discrepancy of physiological signals across various body sites.
Figure 9: Visualization of temporal attention $\alpha_{t}$. The black, blue, green, and red signals represent temporal attention, facial rPPG, acral rPPG, and ground truth PPG signals, respectively.
VI Conclusion
This paper proposes a two-stage deep learning framework consisting of DRP-Net and BBP-Net for estimating heart rate and blood pressure from facial videos. In this paper, we introduce the concept of phase-shifted rPPG signals for analyzing phase discrepancy of pulse waves at facial and acral regions. The DRP-Net extracts facial and acral rPPG signals based on 3D convolution and Siamese-structured heads, and a time domain loss is proposed to supervise the scale of the estimated rPPG signals. The loss function in the frequency domain enabled the learning of rPPG signal phases corresponding to the video sequences of the captured body sites. The BBP-Net estimates SBP and DBP values within a bounded range from the phase-shifted rPPG signals. Phase-shifted rPPG signals provide insights for blood pressure estimation by conveying the time delay of the cardiac cycle. Moreover, the frame interpolation based data augmentation method makes the heart rate distribution wider resulting reduce the bias of the model. Experiments were conducted on the MMSE-HR and V4V databases, and our proposed method achieved superior performance with a significant margin compared to previous methods. Ablation study was thoroughly performed to demonstrate the effectiveness of the phase-shifted rPPG signals for estimating heart rate and blood pressure.
Figure 10: Visualization of spatial attention $\alpha_{s}$. The highlighted areas illustrate regions with higher attention scores.